OpenStudy (abbycross167):

Can someone please help me answer this question? I did a couple steps but I don't know what to do next.... Question~ Solve the systems using the elimination method. y = 2x - 1 y = -x + 7 My Answer (the steps I was able to figure out)~ -2x+y=-1 X+y=7 = -2x+y=-1 Multiply by 1 X+y=7 multiply by 2

OpenStudy (mustbethestars):

2x-1 = -x+7 3x-1 = 7 3x = 8 x =8/3 y= 13,3 once plugging it in Answer (8/3, 13/3)

OpenStudy (abbycross167):

@mustbethestars Thank you very much!!!! I'm sorry for the late reply but my computer messed up